A tennis ball is served from a height of 8 feet and just clears the net that is 3 feet high. The server is 40 feet from the net. If the ball travels in a straight path, how far on the other side of the net does the ball land.
I have sketched this but can't see how to calculate the distance from where the net hits the ground at the 40 ft. to where the ball hits the ground.
Your help would be much appreciated.
Thanks

Answer by Alan3354(31538)

A tennis ball is served from a height of 8 feet and just clears the net that is 3 feet high. The server is 40 feet from the net. If the ball travels in a straight path, how far on the other side of the net does the ball land.
-----------------
8 - 3 = 5 ft (ht above the net)
5/40 = 3/d
5d = 120
d = 24 feet